# A Reduction-Driven Local Search for the Generalized Independent Set Problem

### v1

@@ 4 Reduction Rules
In this section, we develop reduction rules that identify reducible graph structures without compromising optimality. We rely on four operations to reduce the current instance to another one. (1) Edge transformation, which identifies removable edges that can be deleted or transferred to permanent edges; (2) Vertex inclusion, which identifies vertices that are in some set $I\in MGIS(G)$; (3) Vertex exclusion, which identifies vertices that are not in some set $I\in MGIS(G)$. (4) Folding, which identifies vertices that can be merged into a single vertex. Based on the four operations, we will introduce a total of 14 reduction rules.
@@ 4.1 Edge Transformation
If there exists an edge $(u,v)\in E$ such that $p(u,v)=0$, then $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}$ is obtained by removing the edge $(u,v)$ from $G$.
Technical proofs of the correctness of this reduction rule, as well as the following reduction rules, are deferred in Appendix A. R1 implies that removing an edge $(u,v)$ that $p(u,v)=0$ does not change optimality.
@@
If there is an edge $(u,v)$ such that $p(u,v)<M$ and $p(u,v)>min(\tilde{w}(u),\tilde{w}(v))$, then $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}$ is obtained by setting $p(u,v):=M$ in $G$.
Compared to permanent edges, removable edges are more challenging to reduce because of the difficulties of identifying whether both endpoints of a removable edge are in some $S\in MGIS(G)$. R2 allows us to transform removable edges into permanent edges without sacrificing optimality, underpinning many further reduction rules.
@@ 4.2 Neighborhood-based Reductions
If there is a vertex $u\in V$ that $w(u)\geq w^{+}(N(u))$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus N_{p}[u]$, and $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$.
The neighborhood weight reduction generalizes the neighborhood reduction rule in (Lamm et al. 2019) for solving the MIS problem. We further extend R3 by considering more neighboring structures and obtain R4-R6.
@@
If there exists a vertex $u\in V$ such that $w(u)\geq w^{+}(N_{p}(u))+\sum_{x\in N_{r}(u)}\max(0,p(u,x))$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus N_{p}[u]$ and $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$.
@@
If there exists a vertex $u\in V$ such that $w(u)<0$ and $\sum_{v\in N(u)}\min(0,p(u,v))>w(u)$, then $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$.
@@
If there exists a vertex $u\in V$ such that $N_{p}(u)$ is a clique and $w(u)\geq\sum_{v\in N_{r}(u)}\max(0,p(u,v))+\max(0,\max_{v\in N_{p}(u)}\tilde{w}(v))$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus N_{p}[u]$ and $w(v)\coloneqq w(v)-p(u,v)$ for each $v\in N_{r}(u)$.
@@ 4.3.1 Degree-one Vertices Reduction
Case 1: $w(u)\geq p(u,v)$. (1) If $w(u)\geq 0$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G\setminus\{u\}$ and $w(v)\coloneqq w(v)-p(u,v)$. (2) If $w(u)<0$, then $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G\setminus\{u\}$, and $w(v)\coloneqq w(v)-p(u,v)+w(u)$.
Case 2: $w(u)<p(u,v)$. (1) If $w(u)\geq 0$ and $w(u)\geq\tilde{w}(v)$, then $\alpha(G)=\alpha(G\setminus\{u,v\})+w(u)$. (2) If $w(u)\geq 0$ and $w(u)<\tilde{w}(v)$, then $\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G\setminus\{u\}$ and $w(v)\coloneqq w(v)-w(u)$. (3) If $w(u)<0$, then $\alpha(G)=\alpha(G\setminus\{u\})$.
@@ 4.3.2 Degree-two Vertices Reductions
Let $u$ be a vertex with exactly two neighbors $x,y$ and $p(x,y)=M$. The following table shows the reduction rules for reducing $u$ under different conditions.
@@
Case 1: $w(u)\geq\max(p(u,x),p(u,y))$ $w(u)\geq 0$
R8.1.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G\setminus\{u\}$,
$w(x)\coloneqq w(x)-p(u,x)$, $w(y)\coloneqq w(y)-p(u,y)$.
@@
R8.1.2: $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G\setminus\{u\}$,
$w(x)\coloneqq w(x)+w(u)-p(u,x)$, $w(y)\coloneqq w(y)+w(u)-p(u,y)$.
Case 2: $p(u,x)>w(u)\geq p(u,y)$ w.l.o.g, assume $p(u,x)>p(u,y)$
@@
R8.2.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u,x\}$
and $w(y)\coloneqq w(y)-p(u,y)$.
@@
R8.2.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$
$w(x)\coloneqq w(x)-w(u)$, $w(y)\coloneqq w(y)-p(u,y)$.
@@
R8.2.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$
$w(y)\coloneqq w(y)+w(u)-p(u,y)$.
Case 3: $\min(p(u,x),p(u,y))>w(u)$ w.l.o.g, assume $\tilde{w}(x)\geq\tilde{w}(y)$
@@
R8.3.1: $\alpha(G)=\alpha(G\setminus\{u,x,y\})+w(u)$
@@
R8.3.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u,y\}$
$w(x)\coloneqq w(x)-w(u)$.
@@
R8.3.3: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$
$w(x)\coloneqq w(x)-w(u)$, $w(y)\coloneqq w(y)-w(u)$.
$w(u)<0$ R8.3.4: $\alpha(G)=\alpha(G\setminus\{u\})$.
Figure 2: Examples of R8. Suppose that vertex $u$ is adjacent to $x$ and $y$ and, $(x,y)\in E_{p}$. $G_{R}$ represents the residual graph $G\setminus\{u,x,y\}$. $\alpha$ and $\alpha^{\prime}$ indicate the optimal objective value in the original graph and the graph after reduction, respectively.
@@
To describe the $R9$ reduction, we further define $p(u,v)=0$ for $(u,v)\notin E$. Note that this does not alter the objective value in the graph.
@@
Let $u$ be a vertex with two neighbors $x,y$ and $p(x,y)<M$. W.l.o.g, assume $p(u,x)\geq p(u,y)$. The following table summarizes the reduction rules under different conditions.
@@
$w(u)\geq p(u,x)+p(u,y)$
R9.1.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
$w(x)\coloneqq w(x)-p(u,x)$, $w(y)\coloneqq w(y)-p(u,y)$
@@
$w(u)<p(u,x)+p(u,y)$
R9.1.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(x)\coloneqq w(x)-p(u,x)$,
$w(y)\coloneqq w(y)-p(u,y)$, $p(x,y)\coloneqq p(x,y)-p(u,x)-p(u,y)+w(u)$
@@
R9.1.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(x)\coloneqq w(x)+w(u)-p(u,x)$,
$w(y)\coloneqq w(y)+w(u)-p(u,y)$, $p(x,y)\coloneqq p(x,y)+w(u)$.
Case 2: $p(u,x)>w(u)\geq p(u,y)$
@@
$w(u)\geq p(u,x)+p(u,y)$
R9.2.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$, $w(y)\coloneqq w(y)-p(u,y)$,
$w(x)\coloneqq w(x)-w(u)$, $p(x,y)\coloneqq p(x,y)+p(u,x)-w(u)$.
@@
$w(u)<p(u,x)+p(u,y)$
R9.2.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
$w(y)\coloneqq w(y)-p(u,y)$, $w(x)\coloneqq w(x)-w(u)$, $p(x,y)\coloneqq p(x,y)-p(u,y)$.
@@
$w(u)\geq p(u,x)+p(u,y)$
R9.2.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
$w(y)\coloneqq w(y)+w(u)-p(u,y)$, $p(x,y)\coloneqq p(x,y)+p(u,x)$.
@@
$w(u)<p(u,x)+p(u,y)$
R9.2.4: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
$w(y)\coloneqq w(y)+w(u)-p(u,y)$, $p(x,y)\coloneqq p(x,y)+w(u)-p(u,y)$.
Case 3: $p(u,y)>w(u)$ $w(u)\geq 0$
R9.3.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
$w(x)\coloneqq w(x)-w(u)$, $w(y)\coloneqq w(y)-w(u)$, $p(x,y)\coloneqq p(x,y)-w(u)$.
@@
$w(u)<p(u,x)+p(u,y)$
R9.3.2: $\alpha(G)=\alpha(G\setminus\{u\})$.
@@
$w(u)\geq p(u,x)+p(u,y)$
R9.3.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G\setminus\{u\}$,
$p(x,y)\coloneqq p(x,y)-w(u)+p(u,x)+p(u,y)$.
Figure 3: Examples of R9. Suppose that vertex $u$ is adjacent to vertices $x$ and $y$, and $(x,y)\notin E_{p}$.

### v2

@@ 4 Reduction Rules
In this section, we develop reduction rules that identify reducible graph structures without compromising optimality. We rely on four operations to reduce the current instance to a simplified instance. (1) Edge transformation, which identifies removable edges that can be deleted or transformed to permanent edges; (2) Vertex inclusion, which identifies vertices that are in at least one set $I\in MGIS(G)$; (3) Vertex exclusion, which identifies vertices that do not belong to a set $I\in MGIS(G)$. (4) Folding, which identifies vertices that can be merged into a single vertex. Based on the four operations, we introduce a total of 14 reduction rules as follows. For simplicity, we denote the graph obtained after applying a reduction rule by $G^{\prime}=(V^{\prime},E^{\prime}=E_{p}^{\prime}\cup E_{r}^{\prime})$. Unless otherwise specified, we assume that $w^{\prime}(v)$, the profit of a vertex $v\in V^{\prime}$,keeps its original profit $w(v)$ if $v$ also exists in $V$, and $p^{\prime}(e)$, the penalty of a removable edge $e\in E_{r}^{\prime}$, keeps its original penalty $p(e)$ if $e$ also exists in $E_{r}$.
@@ 4.1 Edge Transformation
If there exists an edge $e=\{u,v\}\in E_{r}$ such that $p(e)=0$, then $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}$ is obtained by only removing the edge $e$ from $G$.
The proofs of the correctness of this reduction rule, and the subsequent reduction rules, are deferred to Appendix A. R1 implies removing an edge $e$ with $p(e)=0$ does not change optimality.
@@
If there is an edge $e=\{u,v\}\in E_{r}$ such that $p(e)>min(\tilde{w}(u),\tilde{w}(v))$, then $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}$ is obtained by only moving $e$ from $E_{r}$ to $E_{p}$ in $G$.
Compared to a permanent edge, a removable edge is more challenging to reduce because of the difficulties of identifying whether both endpoints of a removable edge are in a set $S\in MGIS(G)$. R2 allows us to transform a removable edge into a permanent edge without sacrificing optimality, underpinning many further reduction rules.
@@ 4.2 Neighborhood-based Reductions
If there is a vertex $u\in V$ such that $w(u)\geq w^{+}(N(u))$, then $u$ is in at least one set $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus N_{p}[u]]$, and $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$ for each $v\in N_{r}(u)$.
We extend R3 by considering more neighboring structures to obtain R4-R6.
@@
If there exists a vertex $u\in V$ such that $w(u)\geq w^{+}(N_{p}(u))+\sum_{x\in N_{r}(u)}\max(0,p(\{u,x\}))$, then $u$ is in at least one set $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus N_{p}[u]]$, and $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$ for each $v\in N_{r}(u)$.
@@
If there exists a vertex $u\in V$ such that $w(u)<0$ and $\sum_{v\in N_{r}(u)}\min(0,p(\{u,v\}))>w(u)$, then $u$ is not in any $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$.
Next, we show the clique reduction rule. A clique of $G$ is a subset of vertices in which every two vertices are connected by a permanent edge.
@@
If there exists a vertex $u\in V$ such that $N_{p}(u)$ is a clique of $G$ and $w(u)\geq\sum_{v\in N_{r}(u)}\max(0,p(\{u,v\}))+\max(0,\max_{v\in N_{p}(u)}\tilde{w}(v))$, then $u$ is in at least one set $I\in MGIS(G)$, $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus N_{p}[u]]$, and $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$ for each $v\in N_{r}(u)$.
@@ 4.3.1 Degree-one Vertices Reduction
Case 1: $\{u,v\}\in E_{r}$ and $w(u)\geq p(\{u,v\})$. (1) If $w(u)\geq 0$, then $u$ is in at least one set $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G[V\setminus\{u\}]$ and $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})$. (2) If $w(u)<0$, then $\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G[V\setminus\{u\}]$ and $w^{\prime}(v)\coloneqq w(v)-p(\{u,v\})+w(u)$. $u$ is in at least one set $I\in MGIS(G)$ iff $v$ is in at least one set $I\in MGIS(G^{\prime})$.
Case 2: $\{u,v\}\in E_{p}$ or $\{u,v\}\in E_{r}$ but $w(u)<p(\{u,v\})$. (1) If $w(u)\geq 0$ and $w(u)\geq\tilde{w}(v)$, then $u$ is in at least one set $I\in MGIS(G)$, $v\notin I$, and $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u,v\}]$. (2) If $w(u)\geq 0$ and $w(u)<\tilde{w}(v)$, then $u$ and $v$ cannot be both included in the same $I\in MGIS(G)$, we have $\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G[V\setminus\{u\}]$ and $w^{\prime}(v)\coloneqq w(v)-w(u)$. (3) If $w(u)<0$, then $u$ is not in any $I\in MGIS(G)$ and $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$.
@@ 4.3.2 Degree-two Vertices Reductions
For ease of describing $R8$ and $R9$, we assume $p(\{u,v\})=+\infty$ for each $\{u,v\}\in E_{p}$. This assumption does not affect the optimal objective value in the graph.
@@
Let $u$ be a vertex with exactly two neighbors $x,y$ and $\{x,y\}\in E_{p}$. The following table shows the rules for reducing $u$ under different conditions.
@@
Case 1: $w(u)\geq\max(p(\{u,x\}),p(\{u,y\}))$ $w(u)\geq 0$
R8.1.1: $u$ is in at least one set $I\in MGIS(G)$,
$\alpha(G)=\alpha(G^{\prime})+w(u)$, where $G^{\prime}\coloneqq G[V\setminus\{u\}]$,
$w^{\prime}(x)\coloneqq w(x)-p(\{u,x\})$, $w^{\prime}(y)\coloneqq w(y)-p(\{u,y\})$.
@@
R8.1.2: $u$ is in at least one set $I\in MGIS(G)$ iff $x\in I$ or $y\in I$,
$\alpha(G)=\alpha(G^{\prime})$, where $G^{\prime}\coloneqq G[V\setminus\{u\}]$,
$w^{\prime}(x)\coloneqq w(x)+w(u)-p(\{u,x\})$, $w^{\prime}(y)\coloneqq w(y)+w(u)-p(\{u,y\})$.
Case 2: $p(\{u,x\})>w(u)\geq p(\{u,y\})$ w.l.o.g, assume $p(\{u,x\})>p(\{u,y\})$
@@
R8.2.1: $u$ is in at least one set $I\in MGIS(G)$ and $x\notin I$,
$\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u,x\}]$,
and $w^{\prime}(y)\coloneqq w(y)-p(\{u,y\})$.
@@
R8.2.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$,
$w^{\prime}(x)\coloneqq w(x)-w(u)$, $w^{\prime}(y)\coloneqq w(y)-p(\{u,y\})$.
$u$ is in at least one set $I\in MGIS(G)$ iff $x$ is not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
@@
R8.2.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$,
$w^{\prime}(y)\coloneqq w(y)+w(u)-p(\{u,y\})$.
$u$ is in at least one set $I\in MGIS(G)$ iff $y$ is in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
Case 3: $\min(p(\{u,x\}),p(\{u,y\}))>w(u)$ w.l.o.g, assume $\tilde{w}(x)\geq\tilde{w}(y)$
@@
R8.3.1: $u$ is in at least one set $I\in MGIS(G)$,
$\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u,x,y\}]$
@@
R8.3.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u,y\}]$,
$w^{\prime}(x)\coloneqq w(x)-w(u)$.
$u$ is in at least one set $I\in MGIS(G)$ iff $x$ is not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
@@
R8.3.3: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$,
$w^{\prime}(x)\coloneqq w(x)-w(u)$, $w^{\prime}(y)\coloneqq w(y)-w(u)$.
$u$ is in at least one set $I\in MGIS(G)$ iff both $x$ and $y$ are not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
$w(u)<0$
R8.3.4: $u$ is not in any $I\in MGIS(G)$, $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$.
@@
For ease of describing $R9$-$R11$, we further assume $p(\{u,v\})=0$ for $\{u,v\}\notin E$. Note that this does not alter the objective value in the graph. Then, for every pair of vertices $\{u,v\}$, there is a corresponding penalty $p(\{u,v\})$.
@@
Let $u$ be a vertex with two neighbors $x$ and $y$. Assume $p(\{u,x\})\geq p(\{u,y\})$ without loss of generality. If $\{x,y\}\in E_{r}$ or $\{x,y\}\notin E$, the following table summarizes the reduction rules under different conditions.
@@
$w(u)\geq p(\{u,x\})+p(\{u,y\})$
R9.1.1: $u$ is in at least one set $I\in MGIS(G)$,
$\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$,
$w^{\prime}(x)\coloneqq w(x)-p(\{u,x\})$, $w^{\prime}(y)\coloneqq w(y)-p(\{u,y\})$
@@
$w(u)<p(\{u,x\})+p(\{u,y\})$
R9.1.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
$w^{\prime}(x)\coloneqq w(x)-p(\{u,x\})$, $w^{\prime}(y)\coloneqq w(y)-p(\{u,y\})$, $p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})-p(\{u,x\})-p(\{u,y\})+w(u)$
$u$ is in at least one set $I\in MGIS(G)$ iff either $x$ or $y$ is not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
@@
R9.1.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
$w^{\prime}(x)\coloneqq w(x)+w(u)-p(\{u,x\})$, $w^{\prime}(y)\coloneqq w(y)+w(u)-p(\{u,y\})$, $p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})+w(u)$.
$u$ is in at least one set $I\in MGIS(G)$ iff either $x$ or $y$ is in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
Case 2: $p(\{u,x\})>w(u)\geq p(\{u,y\})$
@@
$w(u)\geq p(\{u,x\})+p(\{u,y\})$
R9.2.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
$w^{\prime}(y)\coloneqq w(y)-p(\{u,y\})$, $w^{\prime}(x)\coloneqq w(x)-w(u)$, $p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})+p(\{u,x\})-w(u)$.
$u$ is in at least one set $I\in MGIS(G)$ iff either $x$ is not in or $y$ is in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
@@
$w(u)<p(\{u,x\})+p(\{u,y\})$
R9.2.2: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
$w^{\prime}(y)\coloneqq w(y)-p(\{u,y\})$, $w^{\prime}(x)\coloneqq w(x)-w(u)$, $p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})-p(\{u,y\})$.
$u$ is in at least one set $I\in MGIS(G)$ iff $x$ is not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
@@
$w(u)\geq p(\{u,x\})+p(\{u,y\})$
R9.2.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
$w^{\prime}(y)\coloneqq w(y)+w(u)-p(\{u,y\})$, $p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})+p(\{u,x\})$.
$u$ is in at least one set $I\in MGIS(G)$ iff $y$ is in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
@@
$w(u)<p(\{u,x\})+p(\{u,y\})$
R9.2.4: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
$w^{\prime}(y)\coloneqq w(y)+w(u)-p(\{u,y\})$, $p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})+w(u)-p(\{u,y\})$.
$u$ is in at least one set $I\in MGIS(G)$ iff $x$ is not in and $y$ is in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
Case 3: $p(\{u,y\})>w(u)$ $w(u)\geq 0$
R9.3.1: $\alpha(G)=\alpha(G^{\prime})+w(u)$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
$w^{\prime}(x)\coloneqq w(x)-w(u)$, $w^{\prime}(y)\coloneqq w(y)-w(u)$, $p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})-w(u)$.
$u$ is in at least one set $I\in MGIS(G)$ iff both $x$ and $y$ are not in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
@@
$w(u)<p(\{u,x\})+p(\{u,y\})$
R9.3.2: $u$ is not in any $I\in MGIS(G)$, $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq G[V\setminus\{u\}]$.
@@
$w(u)\geq p(\{u,x\})+p(\{u,y\})$
R9.3.3: $\alpha(G)=\alpha(G^{\prime})$ where $G^{\prime}\coloneqq(V\setminus\{u\},E(G[V\setminus\{u\}])\cup\{\{x,y\}\})$,
$p^{\prime}(\{x,y\})\coloneqq p(\{x,y\})-w(u)+p(\{u,x\})+p(\{u,y\})$.
$u$ is in at least one set $I\in MGIS(G)$ iff both $x$ and $y$ are in at least one set $I^{\prime}\in MGIS(G^{\prime})$.
Figure 2: Examples of $R8$. Suppose that vertex $u$ is adjacent to $x$ and $y$, and $\{x,y\}\in E_{p}$. $G_{R}$ represents the residual graph $G[V\setminus\{u,x,y\}]$. $\alpha$ and $\alpha^{\prime}$ indicate the optimal objective value in the original graph and the graph after reduction, respectively.
Figure 3: Examples of $R9$. Suppose that vertex $u$ is adjacent to vertices $x$ and $y$, and $\{x,y\}\in E_{r}$ or $\{x,y\}\notin E$.